OpenStudy (anonymous):

use the change of base formual to write an expression in terms of logarithms to base e for log5 10.what is the value? (nearest 0.01)

OpenStudy (anonymous):

\[\log_5(10)=\frac{\ln(10)}{\ln(5)}\]

OpenStudy (anonymous):

which comes out to?

OpenStudy (anonymous):

whatever you calculator says

OpenStudy (anonymous):

but i dont have one .__.

OpenStudy (anonymous):

of course you do, your are on a computer

OpenStudy (anonymous):

so whats the answer?